OpenStudy (anonymous):

What is the concentration of NaCl in a solution made by mixing 25.0 g if NaCl with 1550.0 mL water?

OpenStudy (xishem):

Molar concentration of solution constituent, i, is modeled by this equation...\[c_i=\frac{n_i}{V_{sol'n}}\]\[25.0g*\frac{1mol\ NaCl}{58.443g\ NaCl}=0.42777mol\ NaCl\]\[c_i=\frac{0.42777mol}{1.5500L}=0.276M\]